Chem 180 help, calculating heat released , work and delta E

2CH_3OH(l) +3O_2 (g) -> 2CO_2 (g) Delta H = -1452.8 kj/mol

a) what is the heat released by the reaction if 10 mols o_2 reacts with 15 mols CH_3OH?
Answer #1

The given balanced chemical reaction is

      2CH3OH(l) + 3O2(g) --> 2CO2(g), Delta H = -1452.8 kJ/mol

        2 mol            3 mol             2 mol

In the above balanced reaction, stoichiometric molar ratio

= moles of O2(g) / moles of CH3OH(l) = 3 / 2

The given molar ratio = moles of O2(g) / moles of CH3OH(l)

= 10/15 = 2/3, which is less than the stoichiometric molar ratio. Hence O2(g) is the limiting reagent and will be exhausted completely leaving some amount of CH3OH(l) remain unreacted.

Now the amount of CH3OH(l) remain unreacted can be calculated as

3 moles of O2(g) that reacts with the amount of CH3OH(l) = 2 mol

=> 10 moles of O2(g) that will react with the amount of CH3OH(l)

= (2 moles of CH3OH(l) /3 moles of O2(g) ) * 10 moles of O2(g)

= 6.66 mol CH3OH(l)

The heat released when 2 moles of CH3OH(l) reacts = -1452.8 kJ

=> The heat that will release when 6.66 moles of CH3OH(l) reacts

= (-1452.8 kJ)*(6.66 mol CH3OH(l) / 2 mol CH3OH(l) )

= - 4837.8 kJ (answer)